If you can show all the work please, thank you!
What are the units of k in the following rate la
Answer #1

Units of concentration M or mol/L

Units of time seconds (s)

Rate is concentration per unit time,

Both [X] and [Y] are concentration

Rate=k[X][Y]=(M/s), thus

(M/s)=k*M*M

k=\tfrac{1}{M*s}=M^{-1}s^{-1}